OpenStudy (trwatkins1):

Which expressions, when evaluated with x = –9, are negative in value? Choose all answers that are correct. A. –2x – 8 B. –26 – 2x C. 35 + 3x D. –3x + (–40)

OpenStudy (trwatkins1):

@YoungStudier

OpenStudy (trwatkins1):

your the first person i go to for help lol

OpenStudy (youngstudier):

I'm not sure.

OpenStudy (anonymous):

D.

OpenStudy (trwatkins1):

btw its multiple choice

OpenStudy (kendricklamar2014):

A. –2x – 8 = 10 B. –26 – 2x = -8 C. 35 + 3x = 8 D. –3x + (–40) = -13 Answer: Choice B and D @trwatkins1

OpenStudy (trwatkins1):

ok then XD

OpenStudy (kendricklamar2014):

=)